Now the student moves the box up a ramp
(with the same coefficient of friction) inclined
at 10.6
◦ with the horizontal.
b) If the box starts from rest at the bottom
of the ramp and is pulled at an angle of 27.6

with respect to the incline and with the same
183 N force, what is the acceleration up the
ramp?
Answer in units of m/s
2
.
1
Expert's answer
2019-12-24T14:30:26-0500

A good sketch is always a good helper:


Write all forces according to Newton's second law:

"Ox: F\text{ cos}27.6^\circ-\mu_kN-mg\text{ sin}10.6^\circ=ma,\\
Oy: F\text{ sin}27.6^\circ-mg\text{ cos}10.6^\circ+N=0.\\
\space\\
\text{Therefore:}\\
\space\\
N=mg\text{ cos}10.6^\circ-F\text{ sin}27.6^\circ,\\
\space\\
a=\frac{F}{m}\text{ cos}27.6^\circ-\mu_k\frac{N}{m}-g\text{ sin}10.6^\circ=\\
\space\\
=\frac{F}{m}(\text{cos}27.6^\circ+\text{ sin}27.6^\circ)-g(\mu_k\text{ cos}10.6^\circ-\text{ sin}10.6^\circ)=\\
=5.10\text{ m/s}^2."


(The answer is correct if the force is 183 N, the coefficient of kinetic friction is 0.28, and the mass of the box is 41.2 kg).